# Motion of a charged test particle around a static black hole in a monopole magnetic field

### v1

@@ V.3 Magnetic field produced by the electric current caused by charged test particles
The electric current is caused by the motions of charged particles so as to cancel out the background monopole magnetic field. If the electric current is too large, the hovering motions in the plasma lump is highly affected. Thus we consider a lump of charged particles hovering on a black hole and estimate the magnetic field generated by the motion of the charged particles. In order to simplify the calculations, we assume the lump occupies a domain specified by $r_{\rm i}\leq r\leq r_{\rm i}+l$ and $0\leq\theta\leq\Theta$ with
@@
We assume the system is static and axisymmetric, or equivalently, the four-vector potential $A_{a}$ does not depend on $t$ and $\varphi$, and $j^{\alpha}\propto\delta_{3}^{\alpha}$. Then, Maxwell’s equations are given as
@@
$\frac{\partial}{\partial\theta}\left(\sqrt{-g}F^{12}\right)$ $=0,$ (112)
$\frac{\partial}{\partial r}\left(\sqrt{-g}F^{21}\right)$ $=0,$ (113)
@@
$db^{z}=\frac{dIr^{2}\sin^{2}\Theta}{4c^{2}\varepsilon_{0}}\left[r^{2}\sin^{2}\Theta+\left(r_{\rm i}+\frac{l}{2}-r\cos\Theta\right)^{2}\right]^{-3/2}$ (117)
@@
$b^{z}=-\frac{nqr_{\rm g}\sin^{2}\Theta}{8c\varepsilon_{0}{\cal Q}_{\rm M}}\int_{R_{\rm i}}^{R_{\rm i}+L}\frac{R^{4}dR}{(2R-3)}\left[R^{2}-2R\left(R_{\rm i}+\dfrac{L}{2}\right)\cos\Theta+\left(R_{\rm i}+\dfrac{L}{2}\right)^{2}\right]^{-3/2},$ (122)
where $R_{\rm i}=r_{\rm i}/r_{\rm g}$ and $L=l/r_{\rm g}$. In order to estimate the integral in R.H.S. of Eq. (122), we assume $L\ll R_{\rm i}$ and then obtain
$b^{z}=-\frac{nqr_{\rm g}}{8c\varepsilon_{0}{\cal Q}_{\rm M}}\times\frac{R_{\rm i}^{2}}{\sqrt{2}\left(2R_{\rm i}-3\right)}\left[1+{\cal O}\left(\frac{L}{R_{\rm i}}\right)\right]\simeq-\frac{nqr_{\rm g}R_{\rm i}}{16\sqrt{2}\thinspace c\varepsilon_{0}{\cal Q}_{\rm M}},$ (123)
where in the last equality, we use the assumption of $R_{\rm i}\gg 1$. Then, by assuming $R_{\rm i}\gg 1$ and $N\gg 1$, we have
$\left|\frac{b^{z}}{B^{r}|_{r=r_{i}}}\right|$ $\simeq\frac{nqr_{\rm g}R_{\rm i}}{16\sqrt{2}\thinspace c\varepsilon_{0}|{\cal Q}_{\rm M}|}\times\frac{R_{\rm i}^{2}}{N^{2}B_{N}}\simeq\frac{nmR_{\rm i}^{3}}{16\sqrt{2}\thinspace\varepsilon_{0}B_{N}^{2}N^{4}},$ (124)
where we have used Eq. (7) in the last equality. If $|b^{z}/B^{r}|_{r=r_{i}}|\ll 1$, or equivalently,
$n$ $\ll\frac{16\sqrt{2}\thinspace\varepsilon_{0}B_{N}^{2}N^{4}}{mR_{\rm i}^{3}}$
$\simeq 1.20\times 10^{9}\left(\frac{m_{\rm p}}{m}\right)\left(\frac{100r_{\rm g}}{r_{\rm i}}\right)^{3}\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)^{2}\left(\frac{N}{10}\right)^{4}[{\rm m}^{-3}]$ (125)
holds, our analysis can be valid. This condition may be satisfied in a typical HII region.
@@ V.4 Backreaction due to the electromagnetic radiation
$\simeq 2.8\times 10^{5}\left(\frac{100r_{\rm g}}{r}\right)^{2}\left(\frac{N}{10}\right)^{2}\left(\frac{q}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{B^{r}|_{r=Nr_{\rm g}}}{10^{-3}{\rm T}}\right){\rm Hz},$ (126)
@@
By assuming $\left|\dfrac{dr}{dt}\right|\ll\nu r$, we have
@@ VI Summary and discussions
We studied the motion of a charged test particle around a spherically symmetric black hole in a monopole magnetic field. Its radial motion is completely the same as that of the charged test particle in the case with no magnetic field. This fact immediately leads to the same results on the electrification of a black hole as the case with no magnetic field. If plasma composed of non-relativistic protons and electrons whose temperatures are not so different from each other surrounds a spherically symmetric black hole in the monopole magnetic field, selective accretions of protons will occur due to the difference of the masses between a proton and an electron. Since $T_{\rm e}<T_{\rm p}<5T_{\rm e}$ will hold in the situation of our interestMos ; Dexter ; Yuan-Narayan , the black hole will acquire positive charge. The total amount of charge will be very small amount but its effect on the charged particles is very large. For example, the radius of ISCO of an electron can be one order larger than those of the electrically neutral test particlepaper-I . We need to take into account these effects on the plasma surrounding an electrified black hole.
@@
By the same reason, a plasma lump can hover on the black hole in the monopole magnetic field, if the size of the lump is much larger than the Larmor radii of the constituent charged particles $r\vartheta$. Here note that a plasma particle of the Larmor radii $r\vartheta$ has kinetic energy equal to that of the Keplerian motion of the orbital radius $r$. Due to this fact, the mean kinetic energy of plasma particles, or equivalently, the local temperature of the plasma lump may be extremely high. For example, the temperature of protons may be the order of $10^{10}$K when the radial position of the plasma lump is equal to about $100r_{\rm g}$ in the case of the black hole with the mass same as that of SgrA*. Although the temperature of the plasma lump is very high, the thermal radiation will not be emitted since the plasma is too diffuse to be in thermal equilibrium with photons in the situation we consider. It seems to be difficult to observe this high temperature of the plasma, even if it is the case.
It is a crucial prediction that the average velocity of electrons will be equal to that of protons, or in other words, the temperature of electrons is much lower than that of protons, in a plasma lump. This situation is very different from the assumed situation in the discussion of the electrification of a static black hole, where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that bounded lumps of collisionless plasma will not be suppliers of positive electric charge to the black hole. However, as shown in Sec. V.4, electrons in a plasma lump can be selectively accreted on the black hole through the energy loss due to the electromagnetic radiation. Thus, lumps of collisionless plasma bounded around a black hole may be suppliers of negative electric charges to the black hole in a monopole magnetic field.
@@ Appendix A About a test particle initially moving outward
Figure 3: The parameter region in which both Eq. (76) and Eq. (136) are satisfied as a shaded domain in the $(v_{(r)}^{2}$${\cal K}^{2})$-plane.
@@ Appendix B Electric current on the surface of a clump
In the case that charged test particles distribute homogeneously, the electric current exist only on the surface of the clump considered in Sec. V.3. We focus on the electric charge $q$ of test particles are identical to each other and positive, and hence the orbital radii of the charged particles are identical to $l:=r\vartheta$. Then we consider the situation depicted in Fig. 4.
The velocities of particles in 2l<x<0-2l<x<0 are given as
$\bm{v}=\left(-v_{(\varphi)}\sin\psi,-v_{(\varphi)}\cos\psi\right)_{xy}~~~~{\rm with}~~\psi_{-}<\psi<\psi_{+}$ (140)
@@
$\psi_{+}$ $=\frac{3}{2}\pi-\arcsin\left(\frac{x+l}{l}\right).$ (142)
Hence, the average velocity of the particle at $x$ is given as
$\langle\bm{v}\rangle$ $=-\frac{v_{(\varphi)}}{2\pi}\left(\int_{\psi_{-}}^{\psi_{+}}\sin\psi d\psi,\int_{\psi_{-}}^{\psi_{+}}\cos\psi d\psi\right)_{xy}$
$=\frac{v_{(\varphi)}}{2\pi}\left(0,+2\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right]\right)_{xy}$ (143)
@@
$dI=2qndr\int_{-2l}^{0}\frac{v_{(\varphi)}}{2\pi}\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right]dx=\frac{1}{2}qnv_{(\varphi)}ldr=\frac{1}{2}qnv_{(\varphi)}r\vartheta dr,$ (144)
where $n$ is the number density of the charged particle of $q$.

### v2

@@ V.3 Magnetic field produced by the electric current caused by charged test particles
The electric current is caused by the motions of charged particles so as to cancel out the background monopole magnetic field. If the electric current is too large, the hovering motions in the plasma lump is highly affected. Thus we consider a lump of charged particles hovering on a black hole and estimate the magnetic field generated by the motion of the charged particles. In order to simplify the calculations, we assume that the lump occupies a domain specified by $r_{\rm i}\leq r\leq r_{\rm i}+l$ and $0\leq\theta\leq\Theta$ with
@@
We assume that the system is static and axisymmetric, or equivalently, the four-vector potential $A_{a}$ does not depend on $t$ and $\varphi$, and $j^{\alpha}\propto\delta_{3}^{\alpha}$. Then, Maxwell’s equations are given as
@@
$\frac{\partial}{\partial\theta}\left(\sqrt{-g}F^{21}\right)$ $=0,$ (111)
$\frac{\partial}{\partial r}\left(\sqrt{-g}F^{12}\right)$ $=0,$ (112)
@@
$db^{z}=\frac{dIr^{2}\sin^{2}\Theta}{2c^{2}\varepsilon_{0}}\left[r^{2}\sin^{2}\Theta+\left(r_{\rm i}+\frac{l}{2}-r\cos\Theta\right)^{2}\right]^{-3/2}$ (116)
@@
$b^{z}=-\frac{nqr_{\rm g}\sin^{2}\Theta}{4c\varepsilon_{0}{\cal Q}_{\rm M}}\int_{R_{\rm i}}^{R_{\rm i}+L}\frac{R^{4}dR}{(2R-3)}\left[R^{2}-2R\left(R_{\rm i}+\dfrac{L}{2}\right)\cos\Theta+\left(R_{\rm i}+\dfrac{L}{2}\right)^{2}\right]^{-3/2},$ (121)
where $R_{\rm i}=r_{\rm i}/r_{\rm g}$ and $L=l/r_{\rm g}$. By replacing the integration variable $R$ by $\beta$ defined as
$R=R_{\rm i}\left(1+\frac{L}{2R_{\rm i}}\right)\left[\frac{L}{2R_{\rm i}}\sinh\beta+\sqrt{1-\left(\frac{L}{2R_{\rm i}}\right)^{2}}\right],$ (122)
and by assuming $L\ll R_{\rm i}$, we obtain
$b^{z}=-\frac{nqr_{\rm g}}{4c\varepsilon_{0}{\cal Q}_{\rm M}}\times\frac{R_{\rm i}^{2}}{\left(2R_{\rm i}-3\right)}\int^{\ln(\sqrt{2}+1)}_{\ln(\sqrt{2}-1)}\frac{d\beta}{\cosh^{2}\beta}\left[1+{\cal O}\left(\frac{L}{R_{\rm i}}\right)\right]\simeq-\frac{\sqrt{2}\thinspace nqr_{\rm g}R_{\rm i}}{8\thinspace c\varepsilon_{0}{\cal Q}_{\rm M}},$ (123)
where in the last equality, we have used the assumption of $R_{\rm i}\gg 1$. By assuming $R_{\rm i}\gg 1$ and $N\gg 1$, we have
$\left|\frac{b^{z}}{B^{r}|_{r=r_{i}}}\right|$ $\simeq\frac{\sqrt{2}\thinspace nqr_{\rm g}R_{\rm i}}{8\thinspace c\varepsilon_{0}|{\cal Q}_{\rm M}|}\times\frac{R_{\rm i}^{2}}{N^{2}B_{N}}\simeq\frac{\sqrt{2}\thinspace nmR_{\rm i}^{3}}{8\varepsilon_{0}B_{N}^{2}N^{4}},$ (124)
where we have used Eqs. (7) and (38) in the last equality. If $|b^{z}/B^{r}|_{r=r_{i}}|\ll 1$, or equivalently,
$n$ $\ll\frac{4\sqrt{2}\thinspace\varepsilon_{0}B_{N}^{2}N^{4}}{mR_{\rm i}^{3}}$
$\simeq 2.99\times 10^{8}\left(\frac{m_{\rm p}}{m}\right)\left(\frac{100r_{\rm g}}{r_{\rm i}}\right)^{3}\left(\frac{B_{N}}{10^{-3}{\rm T}}\right)^{2}\left(\frac{N}{10}\right)^{4}[{\rm m}^{-3}]$ (125)
holds, our analysis can be valid. This condition is not satisfied in broad-line regions in typical AGNPeterson but a typical HII region.
@@ V.4 Backreaction due to the electromagnetic radiation
$\simeq 2.8\times 10^{5}\left(\frac{100r_{\rm g}}{r}\right)^{2}\left(\frac{N}{10}\right)^{2}\left(\frac{q}{e}\right)\left(\frac{m_{\rm e}}{m}\right)\left(\frac{B_{N}}{10^{-3}{\rm T}}\right){\rm Hz},$ (126)
@@
By assuming that the radial velocity is much smaller compared to the angular velocity, $\left|\dfrac{dr}{dt}\right|\ll\nu r$, we have
@@ VI Summary and discussions
We studied the motion of a charged test particle around a spherically symmetric black hole in a monopole magnetic field. Its radial motion is completely the same as that of the charged test particle in the case with no magnetic field. This fact immediately leads to the same results on the electrification of a black hole as the case with no magnetic field. If plasma composed of non-relativistic protons and electrons whose temperatures are not so different from each other surrounds a spherically symmetric black hole in the monopole magnetic field, selective accretions of protons will occur due to the difference of the masses between a proton and an electron. Since $T_{\rm e}<T_{\rm p}<5T_{\rm e}$ will hold in the situation of our interestMos ; Dexter ; Yuan-Narayan , the black hole will acquire positive charge. The total amount of charge will be a very small amount but its effect on the charged particles is very large. For example, the radius of ISCO of an electron can be one order larger than those of the electrically neutral test particlepaper-I . We need to take into account these effects on the plasma surrounding an electrified black hole.
@@
By the same reason, a plasma lump can hover on the black hole in the monopole magnetic field, if the size of the lump is much larger than the Larmor radii of the constituent charged particles, $r\vartheta$. Here note that a plasma particle of the Larmor radius $r\vartheta$ has kinetic energy equal to that of the Keplerian motion of the orbital radius $r$. Due to this fact, the mean kinetic energy of plasma particles, or equivalently, the local temperature of the plasma lump may be extremely high. For example, the temperature of protons may be the order of $10^{10}$K when the radial position of the plasma lump is equal to about $100r_{\rm g}$ in the case of the black hole with the mass same as that of Sgr A*. Although the temperature of the plasma lump is very high, the thermal radiation will not be emitted since the plasma is too diffuse to be in thermal equilibrium with photons in the situation we consider. It seems to be difficult to observe this high temperature of the plasma, even if it is the case.
It is a crucial prediction that the average velocity of electrons will be equal to that of protons, or in other words, the temperature of electrons is much lower than that of protons, in a plasma lump. This situation is very different from the situation assumed in the discussion of the electrification of a static black hole paper-I , where we have assumed that the velocity distribution of electrons and that of protons are Maxwell one with not so different temperatures from each other. Hence, we should note that bounded lumps of collisionless plasma will not be suppliers of positive electric charge to the black hole. However, as shown in Sec. V.4, electrons in a plasma lump can be selectively accreted on the black hole through the energy loss due to the electromagnetic radiation. Thus, lumps of collisionless plasma bounded around a black hole may be suppliers of negative electric charges to the black hole in a monopole magnetic field.
@@ Appendix A About a test particle initially moving outward
Figure 3: The parameter region in which both Eq. (76) and Eq. (136) are satisfied shown as a shaded domain in the $(v_{(r)}^{2}$${\cal K}^{2})$-plane.
@@ Appendix B Electric current on the surface of a clump
In the case that charged test particles distribute homogeneously in a clump considered in Sec. V.3, the electric current exist only on the surface of the clump. We focus on the situation where the electric charge $q$ of test particles are identical to each other and is positive, and hence the orbital radii of the charged particles are identical to $l:=r\vartheta$. Then we consider the situation depicted in Fig. 4.
The velocities of particles are represented in the form
$v_{(x)}$ $=-v\sin\psi,$ (140)
$v_{(y)}$ $=-v\cos\psi,$ (141)
@@
$v:=\sqrt{v_{(\theta)}^{2}+v_{(\varphi)}^{2}}=\frac{c{\cal K}}{R}$ (142)
[see Eq. (61)].
Assume that the particles are uniformly distributed with the number density $n=n_{0}$ is constant only within the region $x\leq 0$ in their initial state, and that the velocity distribution is isotropic, i.e., the probability with respect to $\psi$ is uniform. Consequently, particles in the domain $2l<x\leq 0$ that satisfy $0\leq\psi<\psi_{-}$ and $\psi_{+}<\psi<2\pi$ oscillate between the domains $x>0$ and $x\leq 0$, where
@@
$\psi_{+}$ $=\frac{3}{2}\pi-\arcsin\left(\frac{x+l}{l}\right)$ (144)
(see Fig. 4). Thus, the number density of the particles in the domain 2l<x<2l-2l<x<2l temporally varies, whereas it in the domain $x\leq-2l$ remain constant at all times. Furthermore, assume that all particles entering the region $x>0$ have been removed for some reason. Then, remaining particles in the domain $2l<x\leq 0$ will have the velocity distribution that is uniform in the domain $\psi_{-}\leq\psi\leq\psi_{+}$ but vanishes outside it, and their number density is given as
$n=n_{0}\times\frac{\psi_{+}-\psi_{-}}{2\pi},$ (145)
The resultant configuration is stationary, and the expectation value of the velocity of each particle in $2l\leq x\leq 0$ is given as
$\langle v_{(x)}\rangle$ $=-\frac{v}{\psi_{+}-\psi_{-}}\int_{\psi_{-}}^{\psi_{+}}\sin\psi d\psi=0,$ (146)
$\langle v_{(y)}\rangle$ $=-\frac{v}{\psi_{+}-\psi_{-}}\int_{\psi_{-}}^{\psi_{+}}\cos\psi d\psi=\frac{2v}{\psi_{+}-\psi_{-}}\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right].$ (147)
@@
$dI$ $=qdr\int_{-2l}^{0}n\langle v_{(y)}\rangle dx$
$=\frac{1}{\pi}qn_{0}vdr\int_{-2l}^{0}\cos\left[\arcsin\left(\frac{x+l}{l}\right)\right]dx$
$=\frac{1}{2}qn_{0}vldr$
$=\frac{1}{2}qn_{0}vr\sin\vartheta dr.$ (148)
From Eq. (51), we have
$vr\sin\vartheta=\frac{cr_{\rm g}{\cal K}^{2}}{\sqrt{{\cal Q}_{\rm M}^{2}+{\cal K}^{2}}}.$ (149)
Because of ${\cal Q}_{\rm M}^{2}\gg{\cal K}^{2}$ in the situation of our interest, we have $vr\sin\vartheta\propto m$ [see Eq. (38)], and hence the contribution of protons to the surface electric current is approximately 2000 times greater than that of electrons in the case of the plasma clump composed of protons and electrons.